# Quantum Bayes’ rule and Petz transpose map from the minimum change principle

### v3

@@ Abstract
where $A_{1}$ and $A_{2}$ are two copies of system $A$ and {|$\{\ket{i}\}$ is a chosen orthonormal basis of ${\mathcal{H}}_{A}$. This pure state has the following marginal states
@@
where the superscript notation $\bullet^{T}$ denotes the transposition taken with respect to the orthonormal basis {|$\{\ket{i}\}$.
@@
Note that the ordering of systems $A$ and $B$ are swapped so that $Q_{\rm rev}\in\mathcal{S}({\mathcal{H}}_{B}\otimes{\mathcal{H}}_{A})$, thus matching Eq. 11. The same representation for the reverse processes has been used for a definition of observational entropy with general priors [45]. In this way, the marginal states become
@@
is the unique solution of the program Eq. 15. Furthermore, if $[\tau,\mathcal{E}(\gamma)]=0$, the above solution coincides with the Petz transpose map of $\mathcal{E}$ with respect to $\gamma$, i.e.,
@@
Figure 2: Example comparing the Petz transpose map with the optimal map obtained from the minimum change principle. The forward map $\mathcal{E}$ models one step of a thermalizing quantum machine where the system interacts with a thermal state $\xi$ [46]. Concretely, we consider qubit systems, and the map is defined as $\mathcal{E}(\rho):=\mathrm{Tr}_{A^{\prime}}[V(\rho\otimes\xi_{A^{\prime}})V^{\dagger}]$, where $V:=\cos\frac{\pi}{8}~\mathds{1}+i\sin\frac{\pi}{8}~U_{\rm sw}$ is the partial swap operator with angle $\frac{\pi}{8}$, $U_{\rm sw}$ is the SWAP gate, and $\xi=0.95|0\rangle\!\!\!\;\langle 0|+0.05|1\rangle\!\!\!\;\langle 1|$. The prior belief is parameterized as $\gamma=(1-p)|0\rangle\!\!\!\;\langle 0|+p|1\rangle\!\!\!\;\langle 1|$, and the reference output state $\tau$ is chosen as $\tau=H\xi H$, where $H$ is the Hadamard gate. The figure plots the value of the objective function of Eq. 15 over $p\in[0.001,0.999]$, for $\mathcal{R}$ chosen to be the Petz transpose map [Eq. 18, orange dashed line] or the optimal map [Eq. 16, blue solid line]. In most cases $[\tau,\mathcal{E}(\gamma)]\neq 0$, and the Petz transpose map is suboptimal; while at $p=1.85-0.9\sqrt{2}\approx 0.577$, where $\mathcal{E}(\gamma)=\mathds{1}/2$ and $[\tau,\mathcal{E}(\gamma)]=0$, the Petz map achieves optimality.
@@
Second, similar to Eq. 5, the optimization contains $\tau$ as a parameter, but the solution Eq. 18 is uniformly optimal regardless of the choice of $\tau$, as long as $[\tau,\mathcal{E}(\gamma)]=0$. The condition $[\tau,\mathcal{E}(\gamma)]=0$, which also guarantees that our solution coincides with the Petz transpose map, is satisfied in some situations of physical interest, such as
@@
Third, when the commutativity condition $[\tau,\mathcal{E}(\gamma)]=0$ is satisfied and the retrodictive channel does not depend on $\tau$, our solution is in line with Jeffrey’s framework of probability kinematics [31], where the update rule relies solely on the prior $\gamma$ and the forward map $\varphi$, independently of the new information $\tau$. However, when $\mathcal{E}(\gamma)$ and $\tau$ do not commute, the quantum case departs from Jeffrey’s framework, and the role of $\tau$ becomes interpretationally significant. If $\tau$ is taken as a subjective prior belief about what the observer expected to observe before obtaining the actual outcome, then its sole purpose is to “guide” the optimization according to the principle of minimum change. In this case, the update rule is given by Eq. 16 with $\sigma$ representing the newly acquired information, and the rule remains independent of $\tau$. In contrast, if $\tau$ is identified with the actual observed evidence—namely, if one sets $\sigma=\tau$ in Eq. 16—then the quantum counterpart of Jeffrey’s retrodicted distribution $\gamma^{\prime}(x)$ from Eq. 2 becomes the quantum state
@@
While the CPTP map $\mathcal{R}$ is always linear, the overall dependence of $\gamma^{\prime}$ on $\tau$ is not, in contrast to the classical case. Again, it is the commutativity condition $[\tau,\mathcal{E}(\gamma)]=0$ that restores the linear dependence of $\gamma^{\prime}$ on $\tau$, and thus in the classical case the two interpretations make no difference. The functional dependence of the retrodicted state on the new information is an important aspect of the problem, which we leave open for future research, especially in view of the axiomatic approach proposed in [25], where linearity of the retrodiction is assumed from the start.
Fourth, when $[\tau,\mathcal{E}(\gamma)]=0$, Theorem 1 can also deal with a non-invertible $\tau$: in this case, we can replace $\tau$ with $\tau_{\epsilon}:=(1-\epsilon)\tau+\epsilon u$, where $\epsilon>0$ and $u$ denotes the uniform (i.e. maximally mixed) state; find the optimal reverse channel, which does not depend on $\tau_{\epsilon}$ as the latter preserves the commutation relation; and finally take the limit for $\epsilon\to 0$. In this way, we see that Theorem 1 does indeed contain the conventional Bayes’ rule (1), where the new information is given as a delta function.
@@
$\Lambda=-\left(\sqrt{\sqrt{\tau}\mathcal{E}(\gamma)\sqrt{\tau}}\right)^{T}$ (21)
@@
Besides fidelity, one may consider other divergences as the figure of merit, such as the Hilbert–Schmidt distance [54] or the Umegaki [55, 56] or Belavkin–Staszewski [57, 58] quantum relative entropies, and wonder whether they lead to the same quantum Bayes’ rule. In the special case of $\tau=\mathcal{E}(\gamma)$, they always do, because the Petz transpose map produces $Q_{\rm fwd}=Q_{\rm ref}$ and any reasonable divergence should reach optimum for two identical operators. However, in the general case $\tau\neq\mathcal{E}(\gamma)$, even if $[\tau,\mathcal{E}(\gamma)]=0$, numerical optimization shows that the optimum points may differ for different divergences. It is thus interesting to explore whether different quantum Bayes’ rules will arise from those various choices.
@@
Acknowledgments.—G. B. and V. S. acknowledge support from the National Research Foundation, Singapore and A*STAR under its CQT Bridging Grant; from the National Research Foundation, Singapore, through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009); and from the Ministry of Education, Singapore, under the Tier 2 grant “Bayesian approach to irreversibility” (Grant No. MOE-T2EP50123-0002). F. B. acknowledges support from MEXT Quantum Leap Flagship Program (MEXT QLEAP) Grant No.  JPMXS0120319794, from MEXT-JSPS Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe,” No. 21H05183, and from JSPS KAKENHI Grant No. 23K03230. G. B. acknowledges support from the Start-up Fund (Grant No. G0101000274) from The Hong Kong University of Science and Technology (Guangzhou).
@@ Appendix A Proof of Theorem 1
$(\sqrt{\tau}\otimes\mathds{1}_{A})\Delta(\sqrt{\tau}\otimes\mathds{1}_{A})+\Lambda^{T}\otimes\mathds{1}_{A}$ =0=0= 0 (42)
$\mathrm{Tr}_{A}[C_{\mathcal{R}}]-\mathds{1}_{B}$ =0=0= 0 (43)
@@
$\Delta=-\tau^{-1/2}\Lambda^{T}\tau^{-1/2}\otimes\mathds{1}_{A}=:\Delta_{B}\otimes\mathds{1}_{A}$ (44)
where $\Delta_{B}:=-\tau^{-1/2}\Lambda^{T}\tau^{-1/2}$. $\Delta_{B}>0$ since $\Delta>0$. By Eq. 43, one gets $\mathds{1}_{B}=\mathrm{Tr}_{A}[C_{\mathcal{R}}^{T}]$. Substituting Eqs. 39 and 44 into this, one gets
@@
$\Lambda^{T}=-\sqrt{\tau}\Delta_{B}\sqrt{\tau}=-\sqrt{\sqrt{\tau}\mathcal{E}(\gamma)\sqrt{\tau}}.$ (52)
@@
Last, if $[\tau,\mathcal{E}(\gamma)]=0$, one has $D=\mathcal{E}(\gamma)^{-1/2}$, and the above equation simplifies to
@@
[…]
@@
[…]

### v4

@@ Abstract
[…]
@@
[…]
@@
Note that the ordering of systems $A$ and $B$ are swapped so that $C_{\mathcal{R}}=\sum_{k,l}|k\rangle\!\!\!\;\langle l|\otimes\mathcal{R}(|k\rangle\!\!\!\;\langle l|)\in\mathcal{L}_{+}({\mathcal{H}}_{B}\otimes{\mathcal{H}}_{A})$ and $Q_{\rm rev}\in\mathcal{S}({\mathcal{H}}_{B}\otimes{\mathcal{H}}_{A})$, thus matching Eq. 11. The same representation for the reverse processes has been used for a definition of observational entropy with general priors [45]. In this way, the marginal states become
@@
is the unique solution of the program Eq. 15. Furthermore, if $\tau,\mathcal{E}(\gamma)]=0$, the above solution coincides with the Petz transpose map of $\mathcal{E}$ with respect to $\gamma$, i.e.,
@@
Figure 2: Example comparing the Petz transpose map with the optimal map obtained from the minimum change principle. The forward map $\mathcal{E}$ models one step of a thermalizing quantum machine where the system interacts with a thermal state $\xi$ [46]. Concretely, we consider qubit systems, and the map is defined as $\mathcal{E}(\rho):=\mathrm{Tr}_{A^{\prime}}[V(\rho\otimes\xi_{A^{\prime}})V^{\dagger}]$, where $V:=\cos\frac{\pi}{8}~\mathds{1}+i\sin\frac{\pi}{8}~U_{\rm sw}$ is the partial swap operator with angle $\frac{\pi}{8}$, $U_{\rm sw}$ is the SWAP gate, and $\xi=0.95|0\rangle\!\!\!\;\langle 0|+0.05|1\rangle\!\!\!\;\langle 1|$. The prior belief is parameterized as $\gamma=(1-p)|0\rangle\!\!\!\;\langle 0|+p|1\rangle\!\!\!\;\langle 1|$, and the reference output state $\tau$ is chosen as $\tau=H\xi H$, where $H$ is the Hadamard gate. The figure plots the value of the objective function of Eq. 15 over $p\in[0.001,0.999]$, for $\mathcal{R}$ chosen to be the Petz transpose map [Eq. 18, orange dashed line] or the optimal map [Eq. 16, blue solid line]. In most cases $\tau,\mathcal{E}(\gamma)]\neq 0$, and the Petz transpose map is suboptimal; while at $p=1.85-0.9\sqrt{2}\approx 0.577$, where $\mathcal{E}(\gamma)=\mathds{1}/2$ and $\tau,\mathcal{E}(\gamma)]=0$, the Petz map achieves optimality.
@@
Second, similar to Eq. 5, the optimization contains $\tau$ as a parameter, but the solution Eq. 18 is uniformly optimal regardless of the choice of $\tau$, as long as $\tau,\mathcal{E}(\gamma)]=0$. The condition $\tau,\mathcal{E}(\gamma)]=0$, which also guarantees that our solution coincides with the Petz transpose map, is satisfied in some situations of physical interest, such as
@@
Third, when the commutativity condition $\tau,\mathcal{E}(\gamma)]=0$ is satisfied and the retrodictive channel does not depend on $\tau$, our solution is in line with Jeffrey’s framework of probability kinematics [31], where the update rule relies solely on the prior $\gamma$ and the forward map $\varphi$, independently of the new information $\tau$. However, when $\mathcal{E}(\gamma)$ and $\tau$ do not commute, the quantum case departs from Jeffrey’s framework, and the role of $\tau$ becomes interpretationally significant. If $\tau$ is taken as a subjective prior belief about what the observer expected to observe before obtaining the actual outcome, then its sole purpose is to “guide” the optimization according to the principle of minimum change. In this case, the update rule is given by Eq. 16 with $\sigma$ representing the newly acquired information, and the rule remains independent of $\tau$. In contrast, if $\tau$ is identified with the actual observed evidence—namely, if one sets $\sigma=\tau$ in Eq. 16—then the quantum counterpart of Jeffrey’s retrodicted distribution $\gamma^{\prime}(x)$ from Eq. 2 becomes the quantum state
@@
While the CPTP map $\mathcal{R}$ is always linear, the overall dependence of $\gamma^{\prime}$ on $\tau$ is not, in contrast to the classical case. Again, it is the commutativity condition $\tau,\mathcal{E}(\gamma)]=0$ that restores the linear dependence of $\gamma^{\prime}$ on $\tau$, and thus in the classical case the two interpretations make no difference. The functional dependence of the retrodicted state on the new information is an important aspect of the problem, which we leave open for future research, especially in view of the axiomatic approach proposed in [25], where linearity of the retrodiction is assumed from the start.
Fourth, when $\tau,\mathcal{E}(\gamma)]=0$, Theorem 1 can also deal with a non-invertible $\tau$: in this case, we can replace $\tau$ with $\tau_{\epsilon}:=(1-\epsilon)\tau+\epsilon u$, where $\epsilon>0$ and $u$ denotes the uniform (i.e. maximally mixed) state; find the optimal reverse channel, which does not depend on $\tau_{\epsilon}$ as the latter preserves the commutation relation; and finally take the limit for $\epsilon\to 0$. In this way, we see that Theorem 1 does indeed contain the conventional Bayes’ rule (1), where the new information is given as a delta function.
@@
$\Lambda=-\frac{1}{2}\left(\sqrt{\sqrt{\tau}\mathcal{E}(\gamma)\sqrt{\tau}}\right)^{T}$ (21)
@@
Besides fidelity, one may consider other divergences as the figure of merit, such as the Hilbert–Schmidt distance [54] or the Umegaki [55, 56] or Belavkin–Staszewski [57, 58] quantum relative entropies, and wonder whether they lead to the same quantum Bayes’ rule. In the special case of $\tau=\mathcal{E}(\gamma)$, they always do, because the Petz transpose map produces $Q_{\rm fwd}=Q_{\rm ref}$ and any reasonable divergence should reach optimum for two identical operators. However, in the general case $\tau\neq\mathcal{E}(\gamma)$, even if $\tau,\mathcal{E}(\gamma)]=0$, numerical optimization shows that the optimum points may differ for different divergences. It is thus interesting to explore whether different quantum Bayes’ rules will arise from those various choices.
@@
Acknowledgments.—The authors would like to thank Maarten Wegewijs for helpful comments. G. B. and V. S. acknowledge support from the National Research Foundation, Singapore and A*STAR under its CQT Bridging Grant; from the National Research Foundation, Singapore, through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009); and from the Ministry of Education, Singapore, under the Tier 2 grant “Bayesian approach to irreversibility” (Grant No. MOE-T2EP50123-0002). F. B. acknowledges support from MEXT Quantum Leap Flagship Program (MEXT QLEAP) Grant No.  JPMXS0120319794, from MEXT-JSPS Grant-in-Aid for Transformative Research Areas (A) “Extreme Universe,” No. 21H05183, and from JSPS KAKENHI Grant No. 23K03230. G. B. acknowledges support from the Start-up Fund (Grant No. G0101000274) from The Hong Kong University of Science and Technology (Guangzhou).
@@ Appendix A Proof of Theorem 1
$\frac{1}{2}(\sqrt{\tau}\otimes\mathds{1}_{A})\Delta(\sqrt{\tau}\otimes\mathds{1}_{A})+\Lambda^{T}\otimes\mathds{1}_{A}$ $=0$ (42)
$\mathrm{Tr}_{A}[C_{\mathcal{R}}]-\mathds{1}_{B}$ $=0$ (43)
@@
$\Delta=-2\tau^{-1/2}\Lambda^{T}\tau^{-1/2}\otimes\mathds{1}_{A}=:\Delta_{B}\otimes\mathds{1}_{A}$ (44)
where $\Delta_{B}:=-2\tau^{-1/2}\Lambda^{T}\tau^{-1/2}$. $\Delta_{B}>0$ since $\Delta>0$. By Eq. 43, one gets $\mathds{1}_{B}=\mathrm{Tr}_{A}[C_{\mathcal{R}}^{T}]$. Substituting Eqs. 39 and 44 into this, one gets
@@
$\Lambda^{T}=-\frac{1}{2}\sqrt{\tau}\Delta_{B}\sqrt{\tau}=-\frac{1}{2}\sqrt{\sqrt{\tau}\mathcal{E}(\gamma)\sqrt{\tau}}.$ (52)
@@
Last, if $\tau,\mathcal{E}(\gamma)]=0$, one has $D=\mathcal{E}(\gamma)^{-1/2}$, and the above equation simplifies to
@@
$\left.\frac{d\Delta}{d\varepsilon}\right|_{\varepsilon=0}=-\int_{0}^{\infty}K(t)HK(t)dt\,.$ (75)
@@
$\left.\frac{d^{2}f}{d\varepsilon^{2}}\right|_{\varepsilon=0}=\frac{1}{2}\mathrm{Tr}\left[\left.\frac{d\Delta}{d\varepsilon}\right|_{\varepsilon=0}H\right]=-\frac{1}{2}\int_{0}^{\infty}\mathrm{Tr}[K(t)HK(t)H]dt\,.$ (76)